OpenStudy (anonymous):

you have 500,000 dollars explain each scenerio and decide which is the best one to use?? Option 1: 6% compounded interest quarterly for 5 years. Option 2: 8% compounded interest annually for 5 years. Option 3: 14.5% simple interest for 10 years.

OpenStudy (anonymous):

\[A = P(1 + \frac{ r }{ n })^{nt}\] P = principal amount r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for n = number of times the interest is compounded per year A = amount of money accumulated after n years, including interest.

OpenStudy (anonymous):

Option 1: 6% compounded interest quarterly for 5 years. P = 500,000 r = (6/100) or 0.06 t = 5 n = 4 So we get: A = 500,000(1 + (0.06/4))^(4*5) A = 673427.50

OpenStudy (anonymous):

Option 2: 8% compounded interest annually for 5 years. P = 500,000 r = (8/100) or 0.08 t = 5 n = 1 So we get: A = 500,000(1 + 0.08))^5 A = 734664.04

OpenStudy (anonymous):

All you really need to do is remember the formula and what every variable in the formula stands for and these problems should be easy from there.